the perimeter of a rectangle is 85ft. the length of the rectangle is 4ft more than its width. what is the equation used to find the width?
Answer by waynest(281)

p = 85
length (L) = w + 4
width = w
p = 2w + 2L
85 = 2w 2(w + 4)
85 = 2w + 2w + 8
85 = 4w + 8
85 - 8 = 4w
77 = 4w
77/4 = 4w/4
19.25 = w
check:
2(19.25) + 2(19.25 + 4) = 85
38.5 + 2(23.25) = 85
38.5 + 46.5 = 85
85 = 85